# Ruelle-Pollicott Decay of Out-of-Time-Order Correlators in Many-Body Systems

### v1

@@ Abstract
The out-of-time-order correlator (OTOC) quantifies information scrambling in quantum systems and serves as a key diagnostic of quantum chaos. In one-body systems with a classical counterpart, the relaxation of the OTOC is governed by Ruelle-Pollicott resonances. For many-body systems lacking a semiclassical limit, recent studies have identified an analogous role played by the Liouvillian spectrum of weakly open extensions of the dynamics, where the slowest decay rate?the Liouvillian gap?encodes relaxation. Here we study the kicked Ising spin chain and show that the long-time exponential decay of the OTOC in the isolated system occurs at a rate equal to twice this intrinsic gap. This correspondence persists even in crossover regimes between integrability and chaos, demonstrating that the Liouvillian spectrum provides a unified framework for understanding relaxation and irreversibility in closed many-body quantum systems.
@@ I Introduction
Understanding relaxation dynamics and the emergence of effective irreversibility in quantum many-body systems is a central problem at the interface of quantum chaos, statistical mechanics, and open-system theory. In classical chaotic dynamics, the intermediate time decay of correlations is governed by the Ruelle-Pollicott (RP) resonances [1, 2, 3] – isolated singularities of the analytically continued spectrum of the Frobenius?Perron operator – which generate effective irreversibility despite the underlying deterministic and reversible motion. In quantum one-body systems with a classical limit, an analogous correspondence is well established through spectral and operator-truncation approaches [4, 5, 6] or through weakly dissipative formulations [7, 8, 9].
@@ V Liouvillian Gap
Parity-resolved analysis – We analyze the Liouvillian spectrum further within each parity subspace. Across the integrable-to-chaotic transition, the gaps of the two subspaces intersect when plotted as a function of the dissipation strength. As a result, the smallest nonzero eigenvalue – and hence the global Liouvillian gap – originates from the even subspace in certain parameter regions, while in others it comes from the odd subspace. This leads to an effective switching between sectors, so that the extracted global Liouvillian gap combines contributions from both subspaces depending on which hosts the slowest decaying mode. Importantly, this does not imply that the Arnoldi-Lindblad method is superior to the approach in Ref. [14]; rather, it allows us to probe parameter regimes?particularly the crossover from integrable to chaotic dynamics?where the behavior of the Liouvillian gaps can be resolved in detail.
@@
Figure 3 compares the Liouvillian gaps computed within each parity subspace?parity-even (blue empty circles) and parity-odd (orange empty squares)?with the global gap obtained via the Arnoldi-Lindblad method (green crosses). Solid lines serve as guides to the eye for the subspace gaps. As the dissipation parameter is varied, the dominant eigenvalue can switch between subspaces: in certain regions the even sector exhibits the smallest eigenvalue, while in others the odd sector becomes dominant. The Arnoldi method naturally captures this transition, since it always selects the eigenvalue with the smallest real part across all subspaces. Moreover, as the system size increases, the spectra of the two parity sectors become increasingly separated, making such crossings less probable. This behavior highlights that the Arnoldi approach consistently identifies the slowest decaying mode of the global dynamics, regardless of the underlying sectoral structure.
@@ VI Relation between OTOC Decay, the Liouvillian Gap and Chaos Parameters
In this section, we bring together the chaos indicators, OTOC decay exponents, and the Liouvillian gap?into a single comparative framework. Previously, we have shown how to compute the normalized level?spacing ratio $\eta$ and participation ratio $\bar{\xi}_{E}$ across different transverse fields $h_{x}$, how to extract the intermediate time exponential decay exponent $\alpha$ of the local correlator $O_{1}^{zz}(l=1,t)$, and how to obtain the Liouvillian gap $\bar{g}$ via quadratic extrapolation in the weak?dissipation limit for $L=10$.
@@
Figure 4(a) displays the chaos indicators $\eta$ (blue upward triangles) and $\bar{\xi}_{E}$ (cyan diamonds) as functions of the transverse field $h_{x}$. Both quantities are computed from the unitary Floquet spectrum at system size $L=12$, and?owing to the symmetries of the model?are analyzed within the parity-even subspace defined with respect to the external reflection operator. These indicators clearly capture the crossover from integrability to chaos.
@@
Overall, this figure highlights that the OTOC decay exponent remains approximately equal to twice the Liouvillian gap for all $h_{x}$ values where scrambling-induced relaxation persists. For $h_{x}>1.00$, the correspondence becomes less pronounced, likely due to finite-size effects?larger system sizes would be required to achieve quantitative convergence of both $\alpha$ and $\bar{g}$ in this regime. Nevertheless, even at these modest sizes, a clear qualitative and quantitative agreement is observed, underscoring the robustness of the relationship between the OTOC decay rate and the Liouvillian gap.
@@ VII Conclusions
A detailed parity-resolved analysis confirms that the intrinsic Liouvillian gap extracted from the Arnoldi?Lindblad method reproduces the global relaxation rate of the system, even when the dominant decay mode shifts between symmetry sectors. This robustness highlights that both the OTOC and the Liouvillian gap respond coherently to changes in the underlying dynamical regime, providing consistent signatures of the transition from integrability to chaos.
The practicality of the Arnoldi?Lindblad approach lies in its ability to access the smallest Liouvillian eigenvalues without constructing the full superoperator, enabling efficient computation of the Liouvillian gap and its extrapolation to the weak-dissipation limit. In contrast to the momentum-resolved analysis of Ref. [13] for systems with periodic boundary conditions, our framework captures the global relaxation properties of the chain under open boundaries, providing a complementary perspective that links spectral relaxation modes directly to information scrambling.
Overall, our results establish a quantitative bridge between unitary and dissipative descriptions of quantum relaxation, extending the Liouvillian?OTOC correspondence beyond the fully chaotic regime. This connection suggests that Liouvillian spectroscopy may serve as a general diagnostic tool for intermediate time dynamics and operator spreading in isolated quantum systems.

### v2

@@ Abstract
The out-of-time-order correlator (OTOC) quantifies information scrambling in quantum systems and serves as a key diagnostic of quantum chaos. In one-body systems with a classical counterpart, the relaxation of the OTOC is governed by Ruelle-Pollicott resonances. For many-body systems lacking a semiclassical limit, recent studies have identified an analogous role played by the Liouvillian spectrum of weakly open extensions of the dynamics, where the slowest decay rate – the Liouvillian gap – encodes relaxation. Here we study the kicked Ising spin chain and show that the long-time exponential decay of the OTOC in the isolated system occurs at a rate equal to twice this intrinsic gap. This correspondence persists even in crossover regimes between integrability and chaos, demonstrating that the Liouvillian spectrum provides a unified framework for understanding relaxation and irreversibility in closed many-body quantum systems.
@@ I Introduction
Understanding relaxation dynamics and the emergence of effective irreversibility in quantum many-body systems is a central problem at the interface of quantum chaos, statistical mechanics, and open-system theory. In classical chaotic dynamics, the intermediate time decay of correlations is governed by the Ruelle-Pollicott (RP) resonances [1, 2, 3] – isolated singularities of the analytically continued spectrum of the Frobenius-Perron operator – which generate effective irreversibility despite the underlying deterministic and reversible motion. In quantum one-body systems with a classical limit, an analogous correspondence is well established through spectral and operator-truncation approaches [4, 5, 6] or through weakly dissipative formulations [7, 8, 9].
@@ V Liouvillian Gap
Parity-resolved analysis – We analyze the Liouvillian spectrum further within each parity subspace. Across the integrable-to-chaotic transition, the gaps of the two subspaces intersect when plotted as a function of the dissipation strength. As a result, the smallest nonzero eigenvalue – and hence the global Liouvillian gap – originates from the even subspace in certain parameter regions, while in others it comes from the odd subspace. This leads to an effective switching between sectors, so that the extracted global Liouvillian gap combines contributions from both subspaces depending on which hosts the slowest decaying mode. Importantly, this does not imply that the Arnoldi-Lindblad method is superior to the approach in Ref. [14]; rather, it allows us to probe parameter regimes – particularly the crossover from integrable to chaotic dynamics – where the behavior of the Liouvillian gaps can be resolved in detail.
@@
Figure 3 compares the Liouvillian gaps computed within each parity subspace – parity-even (blue empty circles) and parity-odd (orange empty squares) – with the global gap obtained via the Arnoldi-Lindblad method (green crosses). Solid lines serve as guides to the eye for the subspace gaps. As the dissipation parameter is varied, the dominant eigenvalue can switch between subspaces: in certain regions the even sector exhibits the smallest eigenvalue, while in others the odd sector becomes dominant. The Arnoldi method naturally captures this transition, since it always selects the eigenvalue with the smallest real part across all subspaces. Moreover, as the system size increases, the spectra of the two parity sectors become increasingly separated, making such crossings less probable. This behavior highlights that the Arnoldi approach consistently identifies the slowest decaying mode of the global dynamics, regardless of the underlying sectoral structure.
@@ VI Relation between OTOC Decay, the Liouvillian Gap and Chaos Parameters
In this section, we bring together the chaos indicators, OTOC decay exponents, and the Liouvillian gap – into a single comparative framework. Previously, we have shown how to compute the normalized level – spacing ratio $\eta$ and participation ratio $\bar{\xi}_{E}$ across different transverse fields $h_{x}$, how to extract the intermediate time exponential decay exponent $\alpha$ of the local correlator $O_{1}^{zz}(l=1,t)$, and how to obtain the Liouvillian gap $\bar{g}$ via quadratic extrapolation in the weak – dissipation limit for $L=10$.
@@
Figure 4(a) displays the chaos indicators $\eta$ (blue upward triangles) and $\bar{\xi}_{E}$ (cyan diamonds) as functions of the transverse field $h_{x}$. Both quantities are computed from the unitary Floquet spectrum at system size $L=12$, and – owing to the symmetries of the model – are analyzed within the parity-even subspace defined with respect to the external reflection operator. These indicators clearly capture the crossover from integrability to chaos.
@@
Overall, this figure highlights that the OTOC decay exponent remains approximately equal to twice the Liouvillian gap for all $h_{x}$ values where scrambling-induced relaxation persists. For $h_{x}>1.00$, the correspondence becomes less pronounced, likely due to finite-size effects – larger system sizes would be required to achieve quantitative convergence of both $\alpha$ and $\bar{g}$ in this regime. Nevertheless, even at these modest sizes, a clear qualitative and quantitative agreement is observed, underscoring the robustness of the relationship between the OTOC decay rate and the Liouvillian gap.
@@ VII Conclusions
A detailed parity-resolved analysis confirms that the intrinsic Liouvillian gap extracted from the Arnoldi-Lindblad method reproduces the global relaxation rate of the system, even when the dominant decay mode shifts between symmetry sectors. This robustness highlights that both the OTOC and the Liouvillian gap respond coherently to changes in the underlying dynamical regime, providing consistent signatures of the transition from integrability to chaos.
The practicality of the Arnoldi-Lindblad approach lies in its ability to access the smallest Liouvillian eigenvalues without constructing the full superoperator, enabling efficient computation of the Liouvillian gap and its extrapolation to the weak-dissipation limit. In contrast to the momentum-resolved analysis of Ref. [13] for systems with periodic boundary conditions, our framework captures the global relaxation properties of the chain under open boundaries, providing a complementary perspective that links spectral relaxation modes directly to information scrambling.
Overall, our results establish a quantitative bridge between unitary and dissipative descriptions of quantum relaxation, extending the Liouvillian-OTOC correspondence beyond the fully chaotic regime. This connection suggests that Liouvillian spectroscopy may serve as a general diagnostic tool for intermediate time dynamics and operator spreading in isolated quantum systems.